OpenStudy (impracticaljoker):

Choose the correct description of the graph of the compound inequality: x − 3 < −9 or x + 5 greater than or equal to 12 (5 points) A number line with a closed circle on −6, shading to the left, and an open circle on 7, shading to the right A number line with an open circle on −6, a closed circle on 7, and shading in between A number line with an open circle on −6, shading to the left, and a closed circle on 7, shading to the right A number line with a closed circle on −6, an open circle on 7, and shading in between

OpenStudy (impracticaljoker):

@quickstudent @TheSmartOne @JoCelM

OpenStudy (quickstudent):

Try typing the equation in this grapher, you will be able to see the graph then and it will be a lot easier.

OpenStudy (quickstudent):

https://www.desmos.com/calculator

OpenStudy (quickstudent):

The link is up there ^^^

OpenStudy (impracticaljoker):

i just insert the equation and it answer it for me?

OpenStudy (jocelm):

x-3<-9 you want to find x, so you need to get rid of the -3, so you add 3 to both sides x-3+3<-9+3 -3+3=0 and -9+3=-6 so x<-6 and then -12+5 on the other so you get x>/=7 So the answer is C.

OpenStudy (impracticaljoker):

idk how to work that

OpenStudy (impracticaljoker):

Thanks JoCel

OpenStudy (jocelm):

Because its less than but not equal to 6 and greather than and equal to 7 so you shade in 7 and do all to the right and you don't shade in -6 and do everything to the left.

OpenStudy (impracticaljoker):

Henry's rocket is designed to launch to a height of 100 feet in the air. The measurements so far have the launch height (h) varying no more than 4 feet. Which inequality below represents this scenario? (5 points) |h − 4| less than or equal to 100 |h + 4| greater than or equal to 100 |h − 100| less than or equal to 4 |h + 100| greater than or equal to 4

OpenStudy (impracticaljoker):

Isn't the answer C?

OpenStudy (impracticaljoker):

I mean D

OpenStudy (jocelm):

No, C. is correct. |h - 100| represents the absolute value of the difference between the actual height and the predicted height of 100 feet. This value varies no more than 4 feet, so |h - 100| <= 4

OpenStudy (impracticaljoker):

Oh